# An Efficient Continual Learning Framework for Multivariate Time Series Prediction Tasks with Application to Vehicle State Estimation

### v1

@@ Abstract
In continual time series analysis using neural networks, catastrophic forgetting (CF) of previously learned models when training on new data domains has always been a significant challenge. This problem is especially challenging in vehicle estimation and control, where new information is sequentially introduced to the model. Unfortunately, existing work on continual learning has not sufficiently addressed the adverse effects of catastrophic forgetting in time series analysis, particularly in multivariate output environments. In this paper, we present EM-ReSeleCT (Efficient Multivariate Representative Selection for Continual Learning in Time Series Tasks), an enhanced approach designed to handle continual learning in multivariate environments. Our approach strategically selects representative subsets from old and historical data and incorporates memory-based continual learning techniques with an improved optimization algorithm to adapt the pre-trained model on new information while preserving previously acquired information. Additionally, we develop a sequence-to-sequence transformer model (autoregressive model) specifically designed for vehicle state estimation. Moreover, we propose an uncertainty quantification framework using conformal prediction to assess the sensitivity of the memory size and to showcase the robustness of the proposed method. Experimental results from tests on an electric Equinox vehicle highlight the superiority of our method in continually learning new information while retaining prior knowledge, outperforming state-of-the-art continual learning methods. Furthermore, EM-ReSeleCT significantly reduces training time, a critical advantage in continual learning applications.
@@ I Introduction
Machine learning models have emerged as powerful tools for addressing complex problems in time series predictions [1]. Recently, Neural networks, due to their non-linear modeling capabilities, have become a popular choice for time series analysis and state estimation tasks [2, 3]. In [4], a novel approach using deep neural networks was proposed to estimate the state of charge (SOC) of electric vehicles (EV). In [5] a hybrid approach was proposed to use neural networks as an aid for model-based approaches in state estimation of dynamical systems with partial information. Recently, there has been significant attention on using Transformer models for time series forecasting tasks, due to their computational efficiency and improved estimation accuracy compared to recurrent neural networks (RNNs) [6, 7].
@@
Unfortunately, these machine learning methods frequently struggle in dynamic environments where the model must continuously adapt to new information. [8]. This issue is particularly significant in time series regression problems, such as robot learning, electric vehicle (EV) battery management, and financial market predictions, where models must adapt to new environments containing new information [9, 10, 11]. The challenge becomes even more critical in vehicle state estimation, where learning-based models must continuously adapt to new data domains encountered by vehicles. Existing work on time series state estimation in dynamic and non-stationary environments has largely focused on local models, which apply local regression fits within each model in each data domain to perform estimations [12, 13, 14]. However, local models face difficulties in noisy environments and can lead to an overwhelming number of models when many domains are introduced, particularly in large datasets. When using neural networks as an alternative, although more powerful, presents challenges when used with static model parameters. In such cases, a pre-trained model struggles to retain previously learned information when exposed to new data, leading to catastrophic forgetting (CF). Unfortunately, retraining the model on both the new and old data domains is impractical due to the high computational complexity involved. Continual learning (CL) aims to enable models to learn from new information while preserving previously acquired knowledge.
@@
Several continual learning approaches have been proposed in the literature, which are generally classified into three groups: regularization-based [8, 15], architectural-based [16], and memory-based [17, 18, 19, 20] methods. In regularization-based methods, important weights or model parameters from previous tasks are preserved to eliminate catastrophic forgetting. Architectural-based methods involve expanding the model’s structure as new tasks are introduced. Memory-based continual learning approaches select a subset of historical data to be retained as a memory, which is then combined with newly received data for training on the current task. The selected subset must effectively represent the old data.
@@
Unfortunately, all these methods have primarily been applied to classification tasks and struggle in addressing the issues related to continual learning in time series regression problems. Recently, ReSeleCT [21] was proposed as a method specifically designed to address catastrophic forgetting in state estimation scenarios. While promising, ReSeleCT has notable limitations. First, it is restricted to univariate cases, focusing on the continual learning and estimation of a single output or state. Second, ReSeleCT employs the exact structure of average gradient of episodic memory (A-GEM) [20] algorithm, which is a memory based CL technique and was originally designed for classification tasks and may not be optimal for regression and prediction problems. Third, the original ReSeleCT implementation relies on recurrent neural networks (RNNs) for state estimation, which may be insufficient for more complex time series problems with substantial data volume and intricate patterns. In such scenarios, more advanced models may be needed, as RNNs can struggle in terms of computational efficiency and representational capability.
@@
In this paper, we introduce our modified scheme on the A-GEM algrorithm, and our approach for multivariate representative or memory selection, to eliminate the adverse effects of catastrophic forgetting in multivariate state estimation. In the experiments, we develop a transformer model-based estimation with an encoder-decoder architecture to improve estimation tasks. Originally introduced in [22], transformer models were designed to enhance the performance of large language models as an alternative to traditional RNN methods. In this work, we adapt transformer models specifically for state estimation tasks, particularly adapted for vehicle datasets. The key contributions in this study are outlined as follows:
@@
  • We apply experimental analysis on real-world data in electric Equinox vehicle state estimation to validate the proposed model, adapting a sequence-to-sequence transformer model for vehicle state estimation with an encoder-decoder mechanism.
@@
The remainder of the paper is organized as follows: Section II presents the problem formulation. In Section III, A-GEM and our refined approach adapted for time series regression problems is introduced. In Section IV, we introduce our approach to continual learning in the multivariate space. Finally, Section V details the experimental analysis of the proposed model in a real-world application on continual learning using a sequence-to-sequence transformer model for vehicle state estimation tasks. Additionally, an uncertainty quantification of the model using the proposed learning method is performed in this section.
@@ II Problem formulation and Background
$\mathbf{y}_{i}=f(\mathbf{x}_{i},H_{i}(D))+\mathbf{\epsilon_{i}},$ (1)
@@
$\ell_{i}(\boldsymbol{\theta})=\|\mathbf{y}_{i}-f(\mathbf{x}_{i}\,;\,\boldsymbol{\theta})\|^{2}.$ (2)
The empirical risk is minimized as follows:
@@
where $|\,D\,|$ represents the total number of data in dataset $D$. Suppose there are $k$ datasets for a given experiment, each comprising time series data structured as in (1):
@@
Here, $n=\sum_{i=1}^{k}n_{i}$ represents the total number of data points in the dataset. In a continual learning setting, datasets from different domains, denoted as $D=D_{1}\cup D_{2}\cup\dots\cup D_{k}$, are not available all at once. Instead, new datasets ($D_{2},\dots,D_{k}$) arrive sequentially, requiring the model to update with new data while retaining previously learned information. In memory-based continual learning, after training on an initial dataset $D_{1}$, the idea is to select a representative subset of $m_{1}$ landmarks, $m_{1}\ll n_{1}$, denoted as $D^{*}_{1}$, to retain in memory. Let us consider the selected landmarks from dataset $D_{1}$ as:
@@ IV Multivariate Representative Selection in time series tasks
The selected memory in the original implementation of A-GEM optimization algorithm is assumed to represent the best subset of the historical data. The original A-GEM, along with our refined version introduced in the previous section use a random subset of historical data as memory points, which is not an optimal selection of past memory. Recently, ReSeleCT [21] was proposed that employs change point detection to identify key locations in the time series data with univariate output. The most informative points are selected as memory points. The selection process involves the Narrowest-Over-Threshold (NOT) method, which applies the Gaussian generalized likelihood ratio (GLR) test to detect change points. The ReSeleCT algorithm iteratively selects representatives from each new dataset and combines them with the existing memory to optimize the neural network using the A-GEM framework. This process allows the model to learn new information while retaining essential knowledge from previous tasks, leading to improved performance in continual learning scenarios. However, a limitation of ReSeleCT is that it is constrained to univariate output time series continual learning scenarios. A generalized approach is required for multivariate cases, where the selection of representative change points must account for the multivariate output environment.
@@
In EM-ReSeleCT, the focus is on identifying specific changes in the output function pattern for each dimension, as these locations are more informative for time series regression tasks and are particularly vulnerable to catastrophic forgetting. To do so, EM-ReSeleCT applies change point detection methods to find specific landmarks based on each output dimension, and then applies an error-based filtering of proximal selected points in each dimension to find the best subset among, and avoid redundant selection. To begin, Let us consider a multivariate space of $p$ input and $q$ output dimensions, as in Eq. 1 where $q$ represents the number of outputs to be estimated. Based on Eq. 4, let the dataset under consideration with $n$ number of instances (including input and output vectors) be defined as:
$D=\left\{\left(\mathbf{x_{1}},\,\mathbf{y_{1}}\right),\,\dots,\,\left(\mathbf{x_{n}},\,\mathbf{y_{n}}\right)\right\}$ (12)
@@
where $\mathbf{x_{i}}$ and $\mathbf{y_{i}}$ are $p$- and $q$-dimensional vectors of inputs and outputs at time instance $i$, respectively:
$\mathbf{x_{i}}=\left\{x_{1i},x_{2i},\dots,x_{pi}\right\}$ (13)
$\mathbf{y_{i}}=\left\{y_{1i},y_{2i},\dots,y_{qi}\right\}$ (14)
To better analyze the dataset over each dimension separately as independent time series dataset, let us define the dataset as input and output dataset pair:
$D=\left\{\mathcal{X},\mathcal{Y}\right\}$ (15)
where, $\mathcal{Y}=\left\{\mathcal{Y}_{1},\mathcal{Y}_{2},\dots,\mathcal{Y}_{q}\right\}$ and $\mathcal{Y}_{i}$ is the time series dataset over dimension $i$ of the output:
$\mathcal{Y}_{i}=\left\{y_{i1},\dots,y_{in}\right\}$ (16)
@@
The objective is to identify representative points from the dataset $D$ that best capture the original dataset’s structure. To achieve this, first a change detection tool called the Narrowest-Over-Threshold (NOT) method [24] is utilized to detect changes in the mean function $\mu_{t}$ of each output dimension of the time series $\mathcal{Y}_{i}$, as $\mathcal{Y}^{\ast}_{i}$ (performing over all dimensions of the output dataset in the historical dataset $\mathcal{Y}$). Therefore, we begin by focusing on selecting a representative subset of the output function for each dimension individually. It is important to note that the choice of the change point detection method depends on both its effectiveness and computational complexity. In this work, we use the NOT algorithm, as it allows for change point selection through linear modeling of the time series dataset (change in the linear mean) while offering excellent computational efficiency. The NOT algorithm begins by partitioning the dataset $\mathcal{Y}_{i}$ into $M$ randomly selected sub-samples. These sub-samples are determined based on observation indices, as depicted in Figure 1, where the interval $(s,e]$ consists of consecutive observations $y_{i(s+1)},\,\dots,\,y_{ie}$. Assuming each sub-sample contains at most one change point, the Gaussian Generalized Likelihood Ratio (GLR) test statistic is computed for potential change points within the interval $(s,e]$. If no change point is present, the likelihood function $L\left(y_{i(s+1)},\,\dots,\,y_{ie};\,\theta\right)$ is maximized to estimate the intercept and slope parameters $\eta=(b,\,m)$ of the mean function $\mu_{t}$ over the range $t\in(s,e]$. Conversely, if a change point occurs at $c\in(s,e]$, the mean function $\mu_{t}$ exhibits distinct intercepts and slopes in the sub-intervals $(s,c]$ and $(c,e]$, characterized by parameters $\eta_{1}$ and $\eta_{2}$, respectively. The GLR test statistic, used to compare the likelihood of no change point versus the presence of a change point at $c$, is formulated as:
@@
$\Re_{\left(s,e\right]}(Y_{i})=\max_{c\in\left\{s+2,\dots,e-2\right\}}\Re_{\left(s,e\right]}^{c}(Y_{i})$ (18)
Thus, within each interval, the change point with the highest likelihood ratio is identified as a potential change point. Once all candidate change points are determined across the $M$ randomly selected intervals, only those with GLR values exceeding a predefined threshold are retained, while the rest are discarded. Finally, among the retained candidates, the change point located within the narrowest interval is selected as the final change point. For further details, refer to [24].
After gathering all change points (landmarks) from each output dimension ($\mathcal{Y}^{\ast}_{1},\mathcal{Y}^{\ast}_{2},\dots,\mathcal{Y}^{\ast}_{q}$), the algorithm collects their corresponding value from the other output dimensions. This results in $q$ distinct sets, each representing the selected landmarks based on one dimension, along with the corresponding outputs from other dimensions:
${\begin{cases}Y^{*}_{1}=\left(\mathbf{y}^{(1)}_{1},\,\dots,\,\mathbf{y}^{(1)}_{(m_{1})}\right)\\ Y^{*}_{2}=\left(\mathbf{y}^{(2)}_{1},\,\dots,\,\mathbf{y}^{(2)}_{(m_{2})}\right)\\ \vdots\\ Y^{*}_{q}=\left(\mathbf{y}^{(q)}_{1},\,\dots,\,\mathbf{y}^{(q)}_{(m_{q})}\right)\end{cases}}$ (19)
where $Y^{*}_{i}$ represents the set of representative points selected based on the $i$th output dimension in the dataset and $\mathbf{y}^{(i)}_{j}=\left\{y^{(i)}_{1j},y^{(i)}_{2j},\dots,y^{(i)}_{qj}\right\}$ is the $j$th captured change point based on the $i$th dimension, along with its corresponding value for other dimensions. one potential solution is to identify change points independently for each output, followed by aggregating these points along with their respective inputs and outputs. However, this approach can result in inefficiencies due to the excessive selection of points. This inefficiency arises from the possibility of redundant point selection, as certain points may already have been chosen for a different output dimension in a neighboring measurement. This problem becomes particularly pronounced in real-world state estimation applications, where the number of measurements is large, and there is minimal variation between consecutive points.
Figure 2 illustrates this issue using the $i$th output time series dataset. In subplot (b), the blue point ($y^{(i)}_{i(k+1)}$) represents the change point detected based on the $i$th time series output dimension, while the red point ($y^{j}_{i(g+1)}$) corresponds to a change point detected in another dimension. As shown in the figure, there is no significant difference between selecting either of these points, and selecting both would be redundant. However, this strategy does not apply to subplot (a), where selecting the red point ($y^{j}_{ig}$) and discarding the blue one ($y^{(i)}_{ik}$) would result in a significant error.
@@
$Y^{*c}=\left\{Y^{*}_{1},Y^{*}_{2},\dots Y^{*}_{q}\right\}$ (20)
@@
Let $Y^{*c}$ represent the set of representative points obtained by concatenating all selected change points for the $j$th dataset. Now, consider the $i$th time series output dimension in $\mathcal{Y}$ within dataset $D$. Based on equation (referenced), the set $\mathcal{Y}^{*}_{i}$ is identified as the set of change points for this dimension. Suppose the $j$th point in this set is denoted by $y_{ij}$.
@@
We assume that within the defined distance $d$, only one change point is detected for dimension $i$. This assumption holds for small distances. To evaluate the significance of this change point, we apply the Generalized Likelihood Ratio (GLR) test. Specifically, we perform the GLR test on the segment from $s_{1}=y_{i(j-\frac{d}{2})}$ (the start of the distance) to the selected point $y*{(i)}_{ij}$, and from this point to $e_{1}=y_{i(j+\frac{d}{2})}$ (the end of the distance):
$\Re_{\left(s_{1},\,e_{1}\right]}^{k(i)}(\mathcal{Y}_{i})=\\ 2\,\log\left[\frac{\sup\limits_{\eta_{1},\eta_{2}}\left\{L\left(y_{i(s_{1}+1)},...,y_{ik};\eta_{1}\right)L\left(y_{i(k+1)},\,\dots,\,y_{ie_{1}};\eta_{2}\right)\right\}}{\sup\limits_{\eta}L\left(y_{i(s_{1}+1)},\,\dots,\,y_{ie_{1}};\eta\right)}\right]$ (21)
@@
Now, for any change point detected in other dimensions, if the corresponding value in the $i$th dimension falls within the defined distance, we apply the GLR test over the same distance as follows:
$\Re_{\left(s_{2},\,e_{2}\right]}^{g(j)}(\mathcal{Y}_{i})=\\ 2\,\log\left[\frac{\sup\limits_{\eta_{1},\eta_{2}}\left\{L\left(y_{i(s_{2}+1)},...,y_{ig};\eta_{1}\right)L\left(y_{i(g+1)},\,\dots,\,y_{ie_{2}};\eta_{2}\right)\right\}}{\sup\limits_{\eta}L\left(y_{i(s_{2}+1)},\,\dots,\,y_{ie_{2}};\eta\right)}\right]$ (22)
@@
Note that the notations for Eqs. (21) and (22) matches the notations in Figure 2 in subplot (a). In other words, we applied GLR test to distance in subplot a, one for point $y^{j}_{ig}$ and one for point $y^{(i)}_{ik}$. We define the error between the two GLR values as:
$e=\Re_{\left(e,\,s\right]}^{k(i)}(\mathcal{Y}_{i})-\Re_{\left(e,\,s\right]}^{g(j)}(\mathcal{Y}_{i})$ (23)
@@
We remove the change point $y^{j}_{ig}$ from the representative set if the error between the two GLR values in Eqs. (21) and (22) is below a predefined threshold. This is because its effect is negligible, and its neighboring point within the proximity can compensate for its absence. Conversely, if the error exceeds the threshold, removing $y^{j}_{ig}$ would significantly impact the representation, so the point is retained in the set.

### v2

@@ Abstract
In continual time series analysis using neural networks, catastrophic forgetting (CF) of previously learned models when training on new data domains has always been a significant challenge. This problem is especially challenging in vehicle estimation and control, where new information is sequentially introduced to the model. Unfortunately, existing work on continual learning has not sufficiently addressed the adverse effects of catastrophic forgetting in time series analysis, particularly in multivariate output environments. In this paper, we present EM-ReSeleCT (Efficient Multivariate Representative Selection for Continual Learning in Time Series Tasks), an enhanced approach designed to handle continual learning in multivariate environments. Our approach strategically selects representative subsets from old and historical data. It incorporates memory-based continual learning techniques with an improved optimization algorithm to adapt the pre-trained model on new information while preserving previously acquired information. Additionally, we develop a sequence-to-sequence transformer model (autoregressive model) specifically designed for vehicle state estimation. Moreover, we propose an uncertainty quantification framework using conformal prediction to assess the sensitivity of the memory size and to showcase the robustness of the proposed method. Experimental results from tests on an electric Equinox vehicle highlight the superiority of our method in continually learning new information while retaining prior knowledge, outperforming state-of-the-art continual learning methods. Furthermore, EM-ReSeleCT significantly reduces training time, a critical advantage in continual learning applications.
@@ I Introduction
Machine learning models have emerged as powerful tools for addressing complex problems in time series predictions [1]. Recently, Neural networks, due to their non-linear modeling capabilities, have become a popular choice for time series analysis and state estimation tasks [2, 3]. In [4], a novel approach using deep neural networks was proposed to estimate the state of charge (SOC) of electric vehicles (EV). In [5], a hybrid approach was proposed to use neural networks to aid model-based approaches in state estimation of dynamical systems with partial information. Recently, significant attention has been paid to using Transformer models for time series forecasting tasks due to their computational efficiency and improved estimation accuracy compared to recurrent neural networks (RNNs) [6, 7].
@@
Unfortunately, these machine learning methods frequently struggle in dynamic environments where the model must continuously adapt to new information. [8]. This issue is particularly significant in time series regression problems, such as robot learning, battery management of electric vehicles (EV), and financial market predictions, where models must adapt to new environments containing new information [9, 10, 11]. The challenge becomes even more critical in vehicle state estimation, where learning-based models must continuously adapt to new data domains encountered by vehicles. Existing work on time series state estimation in dynamic and non-stationary environments has largely focused on local models, which apply local regression fits within each model in each data domain to perform estimations [12, 13, 14]. However, local models face difficulties in noisy environments and can lead to an overwhelming number of models when many domains are introduced, particularly in large datasets. When using neural networks as an alternative, although more powerful, they present challenges when used with static model parameters. In such cases, a pre-trained model struggles to retain previously learned information when exposed to new data, leading to catastrophic forgetting (CF). Unfortunately, retraining the model on both the new and old data domains is impractical due to the high computational complexity involved. Continual learning (CL) aims to enable models to learn from new information while preserving previously acquired knowledge.
@@
Several continual learning approaches have been proposed in the literature, which are generally classified into three groups: regularization-based [8, 15], architectural-based [16], and memory-based [17, 18, 19, 20] methods. In regularization-based methods, important weights or model parameters from previous tasks are preserved to eliminate catastrophic forgetting. Architectural-based methods involve expanding the model’s structure as new tasks are introduced. Memory-based continual learning approaches select a subset of historical data to be retained as memory, which is then combined with newly received data for training on the current task. The selected subset must effectively represent the old data.
@@
Unfortunately, all these methods have primarily been applied to classification tasks and struggle to address the issues related to continual learning in time series regression problems. Recently, ReSeleCT [21] was proposed as a method specifically designed to address catastrophic forgetting in state estimation scenarios. Although promising, ReSeleCT has notable limitations. First, it is restricted to univariate cases, focusing on the continual learning and estimation of a single output or state. Second, ReSeleCT employs the exact structure of the average gradient of episodic memory (A-GEM) algorithm [20], which is a memory-based CL technique and was initially designed for classification tasks and may not be optimal for regression and prediction problems. Third, the original ReSeleCT implementation relies on recurrent neural networks (RNNs) for state estimation, which may be insufficient for more complex time series problems with substantial data volume and intricate patterns. In such scenarios, more advanced models may be needed, as RNNs can struggle in terms of computational efficiency and representational capability.
@@
In this paper, we introduce our modified scheme on the A-GEM algorithm and our approach for multivariate representative or memory selection to eliminate the adverse effects of catastrophic forgetting in multivariate state estimation. In the experiments, we develop a transformer model-based estimation with an encoder-decoder architecture to improve estimation tasks. Originally introduced in [22], transformer models were designed to enhance the performance of large-language models as an alternative to traditional RNN methods. In this work, we adapt transformer models specifically for state estimation tasks, particularly adapted for vehicle datasets. The key contributions in this study are outlined as follows:
@@
  • We apply experimental analysis on real-world data in the electric Equinox vehicle state estimation to validate the proposed model, adapting a sequence-to-sequence transformer model for vehicle state estimation with an encoder-decoder mechanism.
@@
The remainder of the paper is organized as follows: Section II presents the problem formulation. Section III introduces A-GEM and our refined approach adapted for time series regression problems. Section IV introduces our approach to continual learning in the multivariate space. Finally, Section V details the experimental analysis of the proposed model in a real-world application of continual learning using a sequence-to-sequence transformer model for vehicle state estimation tasks. Additionally, this section performs an uncertainty quantification of the model using the proposed learning method.
@@ II Problem formulation and Background
$\mathbf{y}_{i}=f(\mathbf{x}_{i},H_{i}(D))+\boldsymbol{\epsilon}_{i},$ (1)
@@
where $|\,D\,|$ represents the total number of data points in dataset $D$. Suppose there are $k$ datasets for a given experiment, each comprising time series data structured as in (1):
@@
Here, $n=\sum_{i=1}^{k}n_{i}$ represents the total number of data points in the combined dataset. In a continual learning setting, datasets from different domains denoted as $D=D_{1}\cup D_{2}\cup\dots\cup D_{k}$, are not available all at once. Instead, new datasets ($D_{2},\dots,D_{k}$) arrive sequentially, requiring the model to update with new data while retaining previously learned information. In memory-based continual learning, after training on an initial dataset $D_{1}$, the idea is to select a representative subset of $m_{1}$ landmarks, $m_{1}\ll n_{1}$, denoted as $D^{*}_{1}$, to retain in memory. Let us consider the selected landmarks from dataset $D_{1}$ as:
@@ IV Multivariate Representative Selection in time series tasks
The selected memory in the original implementation of the A-GEM optimization algorithm is assumed to represent the best subset of the historical data. The original A-GEM, along with our refined version introduced in the previous section, use a random subset of historical data as memory points, which is not an optimal selection of past memory. Recently, ReSeleCT [21] was proposed, which employs change point detection to identify key locations in the time series data with univariate output. These change points are considered the most informative observations and are selected as memory data. The selection process involves the Narrowest-Over-Threshold (NOT) method, which applies the Gaussian generalized likelihood ratio (GLR) test to detect change points. Employing this change point detection method, the ReSeleCT algorithm iteratively selects representatives from each new dataset (change points) and combines them with the existing memory to optimize the neural network using the A-GEM framework. This process allows the model to learn new information while retaining essential knowledge from previous tasks, leading to improved performance in continual learning scenarios. However, a limitation of ReSeleCT is that it is constrained to univariate output time series continual learning scenarios. A generalized approach is required for multivariate cases, where the selection of representative change points must account for the multivariate output environment.
@@
In EM-ReSeleCT, the focus is on identifying specific changes in the output function pattern for each dimension, as these locations are more informative for time series regression tasks and are particularly vulnerable to catastrophic forgetting. To do so, EM-ReSeleCT applies change point detection methods to find specific landmarks based on each output variable and then applies an error-based filtering of proximal selected points in each output variable to find the best subset and avoid redundant selection. To begin, Let us consider a multivariate setting of $p$ input and $q$ output variables, as in Eq. 1 where $q$ represents the number of outputs to be estimated. Similar to Eq. 3, let $D$ denote the dataset of $n$ instances for the pair of input and output vectors, i.e.
$D=\left\{\left(\mathbf{x}_{1},\,\mathbf{y}_{1}\right),\,\dots,\,\left(\mathbf{x}_{n},\,\mathbf{y}_{n}\right)\right\}\,,$ (11)
@@
where $\mathbf{x}_{j}=(x_{1j},\,\dots,\,x_{pj})^{\prime}$ and $\mathbf{y}_{j}=(y_{1j},\,\dots,\,y_{qj})^{\prime}$ are $p$- and $q$-dimensional vectors of inputs and outputs at time instance $i$, respectively. Additionally, for $i=1,2,\dots,q$, let $\mathcal{Y}_{i}=\left\{y_{i1},\,y_{i2},\dots,\,y_{in}\right\}$ represents the dataset corresponding to the $i_{th}$ coordinate (variable) of the multivariate time series $\mathbf{y}_{1},\,\mathbf{y}_{2},\,\dots,\,\mathbf{y}_{n}$.
@@
The objective is to identify representative points from the dataset $D$ that best capture the original dataset’s structure. To achieve this, first, a change point detection method called the Narrowest-Over-Threshold (NOT) [24] is utilized to detect changes in the mean function $\mu_{t}$ of each output time series $\mathcal{Y}_{i}$. Let us denote the set of the change points by $\mathcal{Y}^{\ast}_{i}$ (performing over all dimensions of the output dataset in the historical dataset $\mathcal{Y}$). Therefore, we begin by focusing on selecting a representative subset of the output data for each output variable individually. It is important to note that the choice of the change point detection method depends on both its effectiveness and computational complexity. In this work, we use the NOT algorithm, as it allows for change point selection through a piece-wise linear modeling of the mean of the time series dataset while offering excellent computational efficiency. The NOT algorithm begins by partitioning the dataset $\mathcal{Y}_{i}$ into $M_{i}$ randomly selected sub-samples. These sub-samples are determined based on observation indices, as depicted in Figure 1, where the interval $(s,e]$ consists of consecutive observations $y_{i(s+1)},\,\dots,\,y_{ie}$. Assuming each sub-sample contains at most one change point, the Gaussian Generalized Likelihood Ratio (GLR) test statistic is computed for potential change points within the interval $(s,e]$. If no change point is present, the likelihood function $L\left(y_{i(s+1)},\,\dots,\,y_{ie};\,\theta\right)$ is maximized to estimate the intercept and slope parameters $\eta=(b,\,m)$ of the linear mean function $\mu_{t}$ over the range $t\in(s,e]$. Conversely, if a change point occurs at $c\in(s,e]$, the mean function $\mu_{t}$ exhibits distinct intercepts and slopes in the sub-intervals $(s,c]$ and $(c,e]$, characterized by parameters $\eta_{1}$ and $\eta_{2}$, respectively. The GLR test statistic, used to compare the likelihood of no change point versus the presence of a change point at $c$, is formulated as:
@@
$\Re_{\left(s,e\right]}(\mathcal{Y}_{i})=\max_{c\in\left\{s+2,\dots,e-2\right\}}\Re_{\left(s,e\right]}^{c}(\mathcal{Y}_{i})$ (13)
Thus, within each interval, the change point with the highest likelihood ratio is identified as a potential change point. Once all candidate change points are determined across the $M_{i}$ randomly selected intervals, only those with GLR values exceeding a predefined threshold are retained, while the rest are discarded. Finally, among the retained candidates, the change point located within the narrowest interval is selected as the final change point. For further details, refer to [24].
After gathering all change points (landmarks) from each output dimension, i.e. $\mathcal{Y}^{\ast}_{1},\mathcal{Y}^{\ast}_{2},\dots,\mathcal{Y}^{\ast}_{q}$, the algorithm collects their corresponding value from the other output dimensions. This results in $q$ distinct sets, each corresponding to the selected change points based on one variable, along with the corresponding outputs from other dimensions:
${\begin{cases}\mathcal{Y}^{**}_{1}=\left(\mathbf{y}^{(1)}_{{}_{1}},\,\dots,\,\mathbf{y}^{(1)}_{{}_{m_{1}}}\right)\\ \mathcal{Y}^{**}_{2}=\left(\mathbf{y}^{(2)}_{{}_{1}},\,\dots,\,\mathbf{y}^{(2)}_{{}_{m_{2}}}\right)\\ \vdots\\ \mathcal{Y}^{**}_{q}=\left(\mathbf{y}^{(q)}_{{}_{1}},\,\dots,\,\mathbf{y}^{(q)}_{{}_{m_{q}}}\right)\end{cases}}$ (14)
where $\mathcal{Y}^{**}_{i}$ represents the set of representative points selected based on the $i$th output variable in the dataset and $\mathbf{y}^{(i)}_{j}=(y^{(i)}_{1j},y^{(i)}_{2j},\dots,y^{(i)}_{qj})$ is an observation vector in the dataset where the mean function of $i^{\rm th}$ coordinate (or output variable) corresponding to $j^{\rm th}$ change when only investigating the $i^{\rm th}$ output variable. One potential solution is to identify change points independently for each output, followed by aggregating these points along with their respective inputs and outputs. However, this approach can result in inefficiencies due to the excessive selection of points. This inefficiency arises from the possibility of redundant point selection, as certain points may already have been chosen for a different output variable in a neighboring measurement. This problem becomes particularly pronounced in real-world state estimation applications, where the number of measurements is large, and there is minimal variation between consecutive points.
Figure 2 illustrates this issue using the $i^{\rm th}$ output time series dataset. In subplot (b), the blue point ($y^{(i)}_{i(k+1)}$) represents the change point detected based on the $i^{\rm th}$ time series output variable, while the red point ($y^{(j)}_{i(g+1)}$) corresponds to a change point detected in another output variable. As shown in the figure, there is no significant difference between selecting either of these points, and selecting both would be redundant. However, this strategy does not apply to subplot (a), where selecting the red point ($y^{(j)}_{ig}$) and discarding the blue one ($y^{(i)}_{ik}$) would result in a significant error.
@@
$\mathcal{Y}^{*c}=\left\{\mathcal{Y}^{**}_{1},\mathcal{Y}^{**}_{2},\dots\mathcal{Y}^{**}_{q}\right\}$ (15)
@@
Let $\mathcal{Y}^{*c}$ represent the set of representative points obtained by concatenating all selected change points for the $j^{\rm th}$ dataset. Now, consider the $i^{\rm th}$ time series output variable in $\mathcal{Y}$ within dataset $D$. Based on equation (16), the set $\mathcal{Y}^{*}_{i}$ is identified as the set of change points for this output variable. Suppose the $j^{\rm th}$ point in this set is denoted by $y_{ij}$.
@@
We assume that within the defined distance $d$, only one change point is detected for the $i^{\rm th}$ output variable. This assumption holds for small distances. To evaluate the significance of this change point, we apply the Generalized Likelihood Ratio (GLR) test. Specifically, we perform the GLR test on the segment from $y_{i(j-\frac{d}{2})}$ (the start of the distance) to the selected point $y^{(i)}_{ij}$, and from this point to $y_{i(j+\frac{d}{2})}$ (the end of the distance):
$\Re_{\left(s_{1},\,e_{1}\right]}^{k,\,i}(\mathcal{Y}_{i})=\\ 2\,\log\left[\frac{\sup\limits_{\eta_{1},\eta_{2}}\left\{L\left(y_{{}_{i\,(s_{1}+1)}},...,y_{{}_{i\,k}};\,\eta_{1}\right)L\left(y_{{}_{i\,(k+1)}},\,\dots,\,y_{{}_{i\,e_{1}}};\eta_{2}\right)\right\}}{\sup\limits_{\eta}L\left(y_{{}_{i\,(s_{1}+1)}},\,\dots,\,y_{{}_{i\,e_{1}}};\eta\right)}\right]$ (16)
@@
where, $s_{1}=j-\frac{d}{2}$ and $e_{1}=j+\frac{d}{2}$, and $y^{i}_{{}_{i\,k}}$ is the $k^{\rm th}$ change point detected based on the $i^{\rm th}$ output variable, and the notation $k,\,i$ in eq. (16) represents the GLR test for $k^{\rm th}$ potential change point for output variable $i$. Now, for any change point detected in other output variables, if the corresponding value in the $i^{\rm th}$ variable falls within the corresponding interval, we apply the GLR test on the same interval as follows:
$\Re_{\left(s_{2},\,e_{2}\right]}^{g,\,j}(\mathcal{Y}_{i})=\\ 2\,\log\left[\frac{\sup\limits_{\eta_{1},\eta_{2}}\left\{L\left(y_{i(s_{2}+1)},...,y_{ig};\eta_{1}\right)L\left(y_{i(g+1)},\,\dots,\,y_{ie_{2}};\eta_{2}\right)\right\}}{\sup\limits_{\eta}L\left(y_{i(s_{2}+1)},\,\dots,\,y_{ie_{2}};\eta\right)}\right]$ (17)
@@
Note that the notations for Eqs. (16) and (17) matches the notations in Figure 2 in subplot (a). In other words, we applied the GLR test to the interval in subplot (a), one for point $y^{j}_{ig}$ and one for point $y^{i}_{ik}$. We define the error between the two GLR values as:
$E=\Re_{\left(e_{1},\,s_{1}\right]}^{k,\,i}(\mathcal{Y}_{i})-\Re_{\left(e_{1},\,s_{1}\right]}^{g,\,j}(\mathcal{Y}_{i})$ (18)
@@
We remove the change point $y^{j}_{ig}$ from the representative set if the difference between the two GLR values in Eqs. (16) and (17) is below a predefined threshold. This is because its effect is negligible, and its neighboring point within the proximity can compensate for its absence. Conversely, if the GLR difference exceeds the threshold, removing $y^{j}_{ig}$ would significantly impact the representation, so the point is retained in the set.